# Distributions and Direct Parametrization for Stable Stochastic State-Space Models

### v1

@@ I Notation
We let $\mathbb{S}^{n}_{>}$ denote the convex cone of positive definite matrices, and $\mathbb{S}^{n}_{\mathrm{skew}}$ the vector space of $n\times n$ skew-symmetric matrices. Let $\operatorname{vec}:\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ be defined by mapping any $X\in\mathbb{R}^{m\times n}$ to its column-stacked vector (with inverse $\operatorname{vec}^{-1}$). Similarly, define $\operatorname{veck}:\mathbb{S}^{n}_{\mathrm{skew}}\to\mathbb{R}^{\frac{n(n-1)}{2}}$ by mapping any skew-symmetric $S$ to the vector of its below-diagonal entries (with inverse $\operatorname{veck}^{-1}$). For a complex vector $v\in\mathbb{C}^{n}$, we let $v^{\dagger}$ denotes its Hermitian (conjugate) transpose. We let $\Gamma$ denote the Gamma function, which generalizes the factorial to complex numbers (with $\Gamma(n)=(n-1)!$ for positive integers $n$). Finally, we let $\mathrm{diag}_{m\times n}(a_{1},\dots,a_{\mathrm{min}\{m,n\}})$ denotes the $m\times n$ rectangular diagonal matrix with $a_{i}$ in the $(i,i)$th entry for $1\leq i\leq\min\{m,n\}$ and zeros elsewhere.
@@ II Introduction
In many real-world applications, one may have strong prior knowledge that the underlying physical or engineered system must satisfy certain stability criteria. When such information is ignored, the identification method may yield models that violate the known stability criteria, thus resulting in poor performance and unreliable long-term predictions. Several approaches have been proposed in the literature to ensure stability or robust performance in the identified models. Constrained optimization strategies have been introduced for deterministic discrete-time systems [10], while specialized regularization techniques have been employed for stochastic discrete-time SSMs [11]. Stability-constrained subspace methods have been discussed in, e.g., [12, 13], and direct parametrization techniques have been developed to guarantee asymptotic stability for deterministic autonomous systems in [13], and quadratic stability in [14] for input-output deterministic and discrete-time linear parameter varying systems. Yet, there remains a scarcity of methods that guarantee stability in the context of continuous-time stochastic SSMs, particularly when dealing with state-dependent and input-dependent process noise.
@@
Overall, our contributions are threefold:
  • We derive a direct parametrization for continuous-time stochastic SSMs with state- and input-dependent noise that inherently satisfies the stochastic bounded-real lemma.
  • We design probabilistic priors over the parametrized model matrices using distributions that enforce stability and discuss sampling techniques from these distributions.
  • We demonstrate, through a simulation example, that the proposed framework yields reliable and stable predictions even when extrapolating beyond the data used for inference.
@@ III Setup
$dx(t)=\left(Ax(t)+Bu(t)\right)dt+\left[Fx(t)~{}Gu(t)\right]dw$ (1a)
$y(t)=Cx(t)+Du(t)$ (1b)
where $A\in\mathbb{R}^{n\times n},F\in\mathbb{R}^{n\times n},G\in\mathbb{R}^{n\times\ell},B\in\mathbb{R}^{n\times\ell},C\in\mathbb{R}^{q\times n},D\in\mathbb{R}^{q\times\ell})$ and $w=\left[w_{1}~{}w_{2}\right]^{\top}$ with $w_{1}$ and $w_{2}$ being real scalar Wiener processes on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ relative to an increasing family $(\mathcal{F}_{t})_{t\geq 0}$ of $\sigma$-algebras $\mathcal{F}_{t}\subset\mathcal{F}$. For the Wiener processes, we assume that
$\mathbb{E}\bigl{[}(w(s)-w(t))(w(s)-w(t))^{\top}\bigr{]}=\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}(s-t),$
with $\rho\in[-1,1]$. We let $\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})$ denote the space of square-integrable $\mathbb{R}^{k}$-valued functions on the probability space $(Ω,\mathcal{F},\mathbb{P})$ i.e., $f\in\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})$ if $\|f\|_{\mathcal{L}^{2}}=\mathbb{E}[\|f\|^{2}_{2}]<\infty$. For $T\in\mathbb{R}_{>}$, we denote by $\mathcal{L}^{2}_{w}\left([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})\right)$ to be the space of nonanticipative stochastic processes $z(\cdot)=\left(z(t)\right)_{t\in[0~{}T]}$ with respect to $\left(\mathcal{F}_{t}\right)_{t\in[0,T]}$ satisfying
[…]
For any $u\in\mathcal{L}^{2}_{w}([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$, and $x_{0}\in\mathbb{R}^{n}$, we denote $x(t;x_{0},u)$ and $y(t;x_{0},u)$ as the solution for (1) with initial condition $x(0)=x_{0}$. The Stochastic Differential Equation (SDE) (1a) with initial condition $x(0)=x_{0}$ and input $u\in\mathcal{L}_{w}([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$ has a unique solution $x(\cdot)=x(\cdot,x_{0},u)\in\mathcal{L}_{w}([0,T],\mathcal{L}^{2}(\Omega,\mathbb{R}^{n}))$ [15]. We now define internal stability for the SSM in (1).
@@
The SSM in (1) is said to be internally stable if there exist constants $c,\alpha>0$ such that the unforced solution of the SDE (1a) $x(\cdot)=x(\cdot,x_{0},0)$ ($u=0$) with $x(0)=x_{0}\in\mathbb{R}^{n}$ satisfies
@@
For $Q\in\mathbb{S}^{n}_{>}$, the SSM in (1) is internally stable if and only if there exists $P\in\mathbb{S}^{n}_{>}$ such that
@@
Note that if $P$ exists then it is unique, and when $Q=\alpha P$, the unforced solution of the SDE (1a) $x(\cdot)=x(\cdot,x_{0},0)$ will satisfy $\mathbb{E}\left[\|x(t)\|^{2}\right]\leq\frac{\lambda_{\mathrm{max}}(P)}{\lambda_{\mathrm{min}}(P)}\exp(-\alpha t)\|x_{0}\|^{2},~{}\text{for all }t>0.$
@@
The SSM (1) is said to be externally stable or $\mathcal{L}^{2}$ input-output stable if, for every $u\in\mathcal{L}_{w}^{2}([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$, we have $y(\cdot)=y(\cdot,0,u)\in\mathcal{L}_{w}^{2}([0~{}T],\mathcal{L}^{2}(\Omega,\mathbb{R}^{q}))$ and there exists a constant $\gamma\geq 0$ such that
$\|y(t)\|_{L_{w}^{2}}\leq\gamma\,\|u(t)\|_{L_{w}^{2}}.$ (4)
If the SSM (1) is externally stable, then we can define an input-output operator $\mathbb{L}:L_{w}^{2}([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))\;\to\;L_{w}^{2}([0~{}T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{q})),$ via $u\mapsto y(\cdot,0,u)$ (see [15]). The induced norm on this operator $\|\mathbb{L}\|:=\sup_{\|u\|_{\mathcal{L}^{2}_{w}}=1}\|y(\cdot,0,u)\|_{\mathcal{L}^{2}_{w}}$ represents the minimal $\gamma>0$ such that (4) is satesfied. This norm is analogous to the $\mathcal{H}_{\infty}$-norm. Therefore, it is often referred to as the stochastic $\mathcal{H}_{\infty}$-norm. In [15], it is shown that if (1) is internally stable, then it is also externally stable and there exists a constant $\gamma>0$ such that $\|\mathbb{L}\|<\gamma$.
@@
    There exists $P\in\mathbb{S}^{n}_{>}$ such that
    $M:=\begin{bmatrix}M_{11}&M_{12}\\ M^{\top}_{12}&M_{22}\end{bmatrix}<0$ (5)
@@ IV-A Distributions over stable $(A,F)$ pairs
For every stable pair $(A,F)$, there exist $P,Q\in\mathbb{S}^{n}_{>}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $\tilde{F}\in\mathbb{R}^{n\times n}$ such that
$A=-\frac{1}{2}P^{-1}(Q+\tilde{F}^{\top}\tilde{F}+S),$ (6a)
$F=L_{P^{-1}}\tilde{F},$ (6b)
where $L_{P^{-1}}$ is the lower-triangular Cholesky factor of $P^{-1}$. Conversely, if $A$ and $F$ are defined by (6) for some $P,Q\in\mathbb{S}^{n}_{>}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $\tilde{F}\in\mathbb{R}^{n\times n}$, then the pair $(A,F)$ are stable.
@@
Let $(A,F)$ be a stable pair, then according to Proposition 1, for a given $Q\in\mathbb{S}^{n}_{>}$, there exists a $P\in\mathbb{S}^{n}_{>}$ such that $A^{\top}P+PA+F^{\top}PF=-Q$. Letting $F:=L_{P^{-1}}\tilde{F}$ for some $\tilde{F}\in\mathbb{R}^{n\times n}$ and $L_{P^{-1}}$ being the lower-triangular Cholesky factor of $P^{-1}$, equation (3) becomes $A^{\top}P+PA+\tilde{F}^{\top}\tilde{F}=-Q$. Letting $\tilde{A}:=PA$, equation (3) becomes $\tilde{A}^{\top}+\tilde{A}=-Q-\tilde{F}^{\top}\tilde{F}$, which implies that $\tilde{A}=-\frac{1}{2}\left(Q+\tilde{F}^{\top}\tilde{F}+S\right)$ which is equivalent to $A=-\frac{1}{2}P^{-1}\left(Q+\tilde{F}^{\top}\tilde{F}+S\right)$ with $S\in\mathbb{S}^{n}_{\mathrm{skew}}$ representing the free skew-symmetric part of $\tilde{A}$. The converse can be shown by a direct substitution of (6) into (3). ∎
It is important to remark that according to Proposition 1, the matrix $Q$ can be fixed to any value. If we fix $Q=\alpha P$, then we can obtain a parametrization which also considers the decay rate $\alpha$ in (2). Additionally, remark that for any eigenvalue $\lambda\in\mathbb{C}$ of $A$ with eigenvector $v\in\mathbb{C}^{n}$, we have
@@
with $\mathrm{Re}(\lambda)=\frac{-v^{\dagger}(Q+\tilde{F}^{\top}\tilde{F})v}{2v^{\dagger}Pv}$, and $\mathrm{Im}(\lambda)=\frac{-v^{\dagger}Sv}{2v^{\dagger}Pv}$. Based on the parametrization in (6), we can define a distribution on stable $(A,F)$ pairs by defining a joint distribution over the parameters $P,\tilde{F},S$, and either fixing $Q\in\mathbb{S}^{n}_{>}$, include it as a random matrix with the joint distribution with the other parameters, or letting $Q=\alpha P$ and including $\alpha$ in the joint distribution for the other parameters. In this paper, we will focus on three possible distributions for $(A,F)$.
@@
Let $P^{-1}\sim\mathrm{Wishart}(k_{p},\Sigma_{p})$ with $k_{p}>n$, $\mathrm{vec}(\tilde{F})\sim\mathcal{N}(\mu_{f},\Sigma_{f})$, and $\mathrm{veck}(S)\sim\mathcal{N}(\mu_{s},\Sigma_{s})$ with $P^{-1},\tilde{F}$ and $S$ being independent. Then we say the pair $(A,F)$ follows
  • The Wishart-Normal-Skew (WNS) distribution if $Q\in\mathbb{S}^{n}_{>}$ is fixed and $(A,F)$ are obtained as in (6).
@@
where $\mathbb{E}\left[P^{-1}\right]=k_{p}\Sigma_{p},~{}\mathbb{E}\left[S\right]=\mathrm{veck}^{-1}(\mu_{s}),~{}\mathbb{E}[\tilde{F}]=\mathrm{vec}^{-1}(\mu_{f}),~{}\mathbb{E}\left[L_{P^{-1}}\right]=L_{\Sigma_{p}}\mathsf{D},$ with $\Sigma_{p}=L_{\Sigma_{p}}L_{\Sigma_{p}}^{\top}$ and $\mathsf{D}$ a diagonal matrix with entries
$\mathsf{D}_{ii}=\sqrt{2}\frac{\Gamma\Bigl{(}\frac{k_{p}-i+2}{2}\Bigr{)}}{\Gamma\Bigl{(}\frac{k_{p}-i+1}{2}\Bigr{)}}.$
@@
$\mathbb{E}[Q]=\begin{cases}Q,&\text{(WSN)},\\ k_{q}\Sigma_{q},&\text{(}Q\text{-WSN)},\\ \mathbb{E}\left[\alpha\right]\mathbb{E}\left[P\right]=k_{\alpha}\theta_{\alpha}k_{P}\Sigma^{-1}_{P},&\text{(}\alpha P\text{-WSN)}.\end{cases}$
Remark 1
One reason to use the Wishart distribution for $P^{-1}\sim\mathrm{Wishart}({k_{p},\Sigma_{p}})$ is the fact that it can be sampled using the Bartlett decomposition $P^{-1}=L_{\Sigma}L_{P^{-1}}L^{\top}_{P^{-1}}L^{\top}_{\Sigma}$, where $L_{\Sigma}$ is the Cholesky factor of $\Sigma_{p}$ and
$L_{P^{-1}}=\begin{bmatrix}c_{1}&0&\cdots&0\\ m_{21}&c_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ m_{n1}&m_{n2}&\cdots&c_{n}\end{bmatrix},$
where $c_{i}^{2}\sim\chi^{2}_{n-i+1}$ and $m_{ij}\sim N(0,1)$ independently. Therefore, we can sample the Cholesky factor $L_{P^{-1}}$ of $P^{-1}$ together without having to sample $P^{-1}$ and then compute a Cholesky factor for it. This approach avoids the computational cost of calculating a Cholesky factor.
The expectation formulas in (8) and the expression for the eigenvalues of $A$ reveal useful insights into how one may choose the parameters for the proposed distributions. Recall that the real part of any eigenvalue $\lambda$ of $A$ in (6) depends on $Q+\tilde{F}^{\top}\tilde{F}$, while the imaginary part is influenced by the skew-symmetric matrix $S$. For instance, when adopting the WNS distribution, one fixes $Q\in\mathbb{S}^{n}_{>}$ and draws samples for $P^{-1}$, $\tilde{F}$, and $S$. This approach ties the mean-squared decay properties (through $Q$) to a known baseline while permitting flexibility in both the real part of the eigenvalues (through the distribution of $P$ and $\tilde{F}$) and their oscillatory behavior (through $S$). Suppose one wants to quantify the uncertainty in the mean-squared decay properties with the WNS. In that case, it is possible by either modifying the uncertainty in $P^{-1}$ through $\Sigma_{p}$ and $k_{p}$ or modifying the uncertainty in $\tilde{F}$ through $\Sigma_{f}$. However, this will also alter the uncertainty in the drift part’s oscillatory behavior or the diffusion part’s uncertainty. In contrast, the $Q$-WNS distribution treats $Q$ itself as random, thereby enabling us to quantify the uncertainty in mean-square decay properties without affecting the uncertainty in the oscillatory properties or the diffusion term.

### v2

@@ I Notation
We let $\mathbb{R}_{>0}=(0,\infty)$, $\mathbb{R}_{\geq 0}=[0,\infty)$, $\mathbb{S}^{n}_{>0}$ denote the convex cone of positive definite matrices, and $\mathbb{S}^{n}_{\mathrm{skew}}$ the vector space of $n\times n$ skew-symmetric matrices. Let $\operatorname{vec}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ be defined by mapping any $X\in\mathbb{R}^{m\times n}$ to its column-stacked vector (with inverse $\operatorname{vec}^{-1}$). Similarly, define $\operatorname{veck}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{n}_{\mathrm{skew}}\to\mathbb{R}^{\frac{n(n-1)}{2}}$ by mapping any skew-symmetric $S$ to the vector of its below-diagonal entries (with inverse $\operatorname{veck}^{-1}$). For a complex vector $v\in\mathbb{C}^{n}$, we let $v^{\dagger}$ denotes its Hermitian (conjugate) transpose. We let $\Gamma$ denote the Gamma function, which generalizes the factorial to complex numbers (with $\Gamma(n)=(n-1)!$ for positive integers $n$). Finally, we let $\mathrm{diag}_{m\times n}(a_{1},\dots,a_{\mathrm{min}\{m,n\}})$ denotes the $m\times n$ rectangular diagonal matrix with $a_{i}$ in the $(i,i)$th entry for $1\leq i\leq\min\{m,n\}$ and zeros elsewhere.
@@ II Introduction
In many real-world applications, one may have strong prior knowledge that the underlying physical or engineered system must satisfy certain stability criteria. When such information is ignored, the identification method may yield models that violate the known stability criteria, thus resulting in poor performance and unreliable long-term predictions. Several approaches have been proposed in the literature to ensure stability or robust performance in the identified models. Constrained optimization strategies have been introduced for deterministic discrete-time systems [10], while specialized regularization techniques have been employed for stochastic discrete-time SSMs [11]. Stability-constrained subspace methods have been discussed in, e.g., [12, 13], while kernel-based regularization techniques guarantee stability for deterministic continuous- and discrete-time systems in [14]. Direct parametrization approaches have been developed to ensure asymptotic stability in deterministic autonomous systems [13] and quadratic stability in input–output deterministic discrete-time linear parameter-varying systems [15]. Yet, methods ensuring stability for continuous-time stochastic SSMs with state and input-dependent (multiplicative) noise are scarce despite their prevalence in diverse applications [16, 17] and the control literature [18].
@@
Overall, our contributions are threefold: (i) We derive a direct parametrization for continuous‑time stochastic SSMs with state‑ and input‑dependent noise that inherently satisfies the stochastic bounded‑real lemma; (ii) We design probabilistic priors over the parametrized model matrices using distributions that enforce stability and discuss sampling techniques from these distributions; and (iii) We demonstrate, through a simulation example, that the proposed framework yields reliable and stable predictions even when extrapolating beyond the data used for inference.
@@ III Setup
$dx(t)=\left(Ax(t)+Bu(t)\right)dt+\left[Fx(t)~Gu(t)\right]dw(t),$ (1a)
$y(t)=Cx(t)+Du(t),$ (1b)
where $A\in\mathbb{R}^{n\times n},F\in\mathbb{R}^{n\times n},G\in\mathbb{R}^{n\times\ell},B\in\mathbb{R}^{n\times\ell},C\in\mathbb{R}^{q\times n},D\in\mathbb{R}^{q\times\ell})$ and $w=\left[w_{1}~w_{2}\right]^{\top}$ with $w_{1}$ and $w_{2}$ being real scalar Wiener processes on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ relative to an increasing family $(\mathcal{F}_{t})_{t\geq 0}$ of $\sigma$-algebras $\mathcal{F}_{t}\subset\mathcal{F}$. All stochastic integrals are understood in the Itô sense. For the Wiener processes, we assume that $\mathbb{E}\bigl[(w(s)-w(t))(w(s)-w(t))^{\top}\bigr]=\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}(s-t),$ with $\rho\in[-1,1]$. We let $\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})$ denote the space of square-integrable $\mathbb{R}^{k}$-valued functions on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ i.e., $f\in\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})$ if $\|f\|_{\mathcal{L}^{2}}=\mathbb{E}[{f^{\top}f}]<\infty$. For $T\in\mathbb{R}_{>0}$, we denote by $\mathcal{L}^{2}_{w}\left([0~T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{k})\right)$ the space of nonanticipative stochastic processes $z(\cdot)=\left(z(t)\right)_{t\in[0~T]}$ with respect to $\left(\mathcal{F}_{t}\right)_{t\in[0,T]}$ satisfying
$\|z\|^{2}_{\mathcal{L}^{2}_{w}}=\int^{T}_{0}\|z(t)\|_{\mathcal{L}^{2}}dt<\infty.$
For any $u\in\mathcal{L}^{2}_{w}([0~T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$, and $x_{0}\in\mathbb{R}^{n}$, we denote $x(t;x_{0},u)$ and $y(t;x_{0},u)$ as the solution for (1) with initial condition $x(0)=x_{0}$. The Stochastic Differential Equation (SDE) (1a) with initial condition $x(0)=x_{0}$ and input $u\in\mathcal{L}_{w}([0~T];\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$ has a unique solution $x(\cdot)=x(\cdot;x_{0},u)\in\mathcal{L}_{w}([0,T],\mathcal{L}^{2}(\Omega,\mathbb{R}^{n}))$ [19]. We now define internal stability for the SSM in (1).
@@
The SSM in (1) is said to be internally stable if there exist constants $c,\alpha>0$ such that the unforced solution of the SDE (1a) $x(\cdot)=x(\cdot;x_{0},0)$ ($u=0$) with $x(0)=x_{0}\in\mathbb{R}^{n}$ satisfies
@@
For $Q\in\mathbb{S}^{n}_{>0}$, the SSM in (1) is internally stable if and only if there exists $P\in\mathbb{S}^{n}_{>0}$ such that
@@
Note that if $P$ exists then it is unique, and when $Q=\alpha P$, the unforced solution of the SDE (1a) $x(\cdot)=x(\cdot;x_{0},0)$ will satisfy $\mathbb{E}\left[\|x(t)\|^{2}\right]\leq\frac{\lambda_{\mathrm{max}}(P)}{\lambda_{\mathrm{min}}(P)}\exp(-\alpha t)\|x_{0}\|^{2},~\text{for all }t>0.$
@@
The SSM (1) is said to be externally stable or $\mathcal{L}^{2}$ input-output stable if, for every $u\in\mathcal{L}_{w}^{2}({\mathbb{R}_{\geq 0}};\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))$, we have $y(\cdot)=y(\cdot;0,u)\in\mathcal{L}_{w}^{2}({\mathbb{R}_{\geq 0}};\mathcal{L}^{2}(\Omega,\mathbb{R}^{q}))$ and there exists a constant $\gamma\geq 0$ such that
$\|y(t)\|_{\mathcal{L}_{w}^{2}}\leq\gamma\,\|u(t)\|_{\mathcal{L}_{w}^{2}}.$ (4)
If the SSM (1) is externally stable, then we can define an input-output operator $\mathbb{L}\mathrel{\mathop{\ordinarycolon}}\mathcal{L}_{w}^{2}({\mathbb{R}_{\geq 0}};\mathcal{L}^{2}(\Omega,\mathbb{R}^{\ell}))\!\to\!\mathcal{L}_{w}^{2}({\mathbb{R}_{\geq 0}};\mathcal{L}^{2}(\Omega,\mathbb{R}^{q})),$ via $u\!\mapsto\!y(\cdot;0,u)$ (see [19]). The induced norm on this operator $\|\mathbb{L}\|\mathrel{\mathop{\ordinarycolon}}=\sup_{\|u\|_{\mathcal{L}^{2}_{w}}=1}\|y(\cdot;0,u)\|_{\mathcal{L}^{2}_{w}}$ represents the minimal $\gamma>0$ such that (4) is satisfied. This norm is analogous to the $\mathcal{H}_{\infty}$-norm. Therefore, it is often referred to as the stochastic $\mathcal{H}_{\infty}$-norm. In [19], it is shown that if (1) is internally stable, then it is also externally stable and there exists a constant $\gamma>0$ such that $\|\mathbb{L}\|<\gamma$.
@@
    There exists $P\in\mathbb{S}^{n}_{>0}$ such that
    $M\mathrel{\mathop{\ordinarycolon}}=\begin{bmatrix}M_{11}&M_{12}\\ M^{\top}_{12}&M_{22}\end{bmatrix}<0$ (5)
@@ IV-A Distributions over stable $(A,F)$ pairs
For every stable pair $(A,F)$, there exist $P,Q\in\mathbb{S}^{n}_{>0}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $\tilde{F}\in\mathbb{R}^{n\times n}$ such that
$A=-\frac{1}{2}P^{-1}(Q+\tilde{F}^{\top}\tilde{F}+S),~F=L_{P^{-1}}\tilde{F}$, (6)
where $L_{P^{-1}}$ is the lower-triangular Cholesky factor of $P^{-1}$. Conversely, if $A$ and $F$ are defined by (6) for some $P,Q\in\mathbb{S}^{n}_{>0}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $\tilde{F}\in\mathbb{R}^{n\times n}$, then the pair $(A,F)$ are stable.
@@
Let $(A,F)$ be a stable pair, then according to Proposition 1, for any given $Q\in\mathbb{S}^{n}_{>0}$, there exists a $P\in\mathbb{S}^{n}_{>0}$ such that $Q=A^{\top}P+PA+F^{\top}PF=\tilde{A}^{\top}+\tilde{A}+\tilde{F}^{\top}\tilde{F}$ where the second equality follows by defining $\tilde{A}\mathrel{\mathop{\ordinarycolon}}=PA$ and $\tilde{F}\mathrel{\mathop{\ordinarycolon}}=L_{P^{-1}}^{-1}F$. With $S\in\mathbb{S}^{n}_{\mathrm{skew}}$ representing the free skew-symmetric part of $\tilde{A}$ it follows that $\tilde{A}=-\frac{1}{2}\left(Q+\tilde{F}^{\top}\tilde{F}+S\right)$ and therefore $A=-\frac{1}{2}P^{-1}\left(Q+\tilde{F}^{\top}\tilde{F}+S\right)$. The converse can be shown by a direct substitution of (6) into (3). ∎
It is important to remark that according to Proposition 1, the matrix $Q$ can be fixed to any value. If we fix $Q=\alpha P$, then we can obtain a parametrization which also considers the decay rate $\alpha$ in (2). Additionally, note that for any eigenvalue $\lambda\in\mathbb{C}$ of $A$ with eigenvector $v\in\mathbb{C}^{n}$, we have $\frac{1}{2}P^{-1}(Q+\tilde{F}^{\top}\tilde{F}+S)v=\lambda v\Rightarrow\frac{-1}{2}(Q+\tilde{F}^{\top}\tilde{F}+S)v=\lambda Pv$, which means
@@
with $\mathrm{Re}(\lambda)\!=\!\frac{-v^{\dagger}(Q+\tilde{F}^{\top}\tilde{F})v}{2v^{\dagger}Pv}$, and $\mathrm{Im}(\lambda)\!=\!\frac{-v^{\dagger}Sv}{2v^{\dagger}Pv}$. Based on the parametrization in (6), we can define a distribution on stable $(A,F)$ pairs by defining a joint distribution over the parameters $P,\tilde{F},S$, and either fixing $Q\in\mathbb{S}^{n}_{>0}$, include it as a random matrix with the joint distribution with the other parameters, or letting $Q=\alpha P$ and including $\alpha$ in the joint distribution for the other parameters. In this paper, we will focus on three possible distributions for $(A,F)$.
@@
Let $P^{-1}\sim\mathrm{Wishart}(k_{p},\Sigma_{p})$ with $k_{p}\geq n$, $\mathrm{vec}(\tilde{F})\sim\mathcal{N}(\mu_{f},\Sigma_{f})$, and $\mathrm{veck}(S)\sim\mathcal{N}(\mu_{s},\Sigma_{s})$ with $P^{-1},\tilde{F}$ and $S$ being independent. Then we say the pair $(A,F)$ follows
  • The Wishart-Normal-Skew (WNS) distribution if $Q\in\mathbb{S}^{n}_{>0}$ is fixed and $(A,F)$ are obtained as in (6).
@@
Remark 1
One reason to use the Wishart distribution for $P^{-1}\sim\mathrm{Wishart}({k_{p},\Sigma_{p}})$ is the fact that it can be sampled using the Bartlett decomposition $P^{-1}=L_{\Sigma_{p}}\tilde{L}_{P^{-1}}\tilde{L}^{\top}_{P^{-1}}L^{\top}_{\Sigma_{p}}$, where $L_{\Sigma_{p}}$ is the Cholesky factor of $\Sigma_{p}$ and
$\tilde{L}_{P^{-1}}=\begin{bmatrix}c_{1}&0&\cdots&0\\ m_{21}&c_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ m_{n1}&m_{n2}&\cdots&c_{n},\end{bmatrix}$ (8)
where $c_{i}^{2}\sim\chi^{2}_{n-i+1}$ and $m_{ij}\sim N(0,1)$ independently [20]. Therefore, we can sample the Cholesky factor $L_{P^{-1}}=L_{\Sigma_{p}}\tilde{L}_{P^{-1}}$ of $P^{-1}$ without having to sample $P^{-1}$ and then compute a Cholesky factor for it. This approach avoids the computational cost of calculating a Cholesky factor.
@@
where $\mathbb{E}\left[P^{-1}\right]\!=\!k_{p}\Sigma_{p},~\mathbb{E}\left[S\right]\!=\!\mathrm{veck}^{-1}(\mu_{s}),~\mathbb{E}[\tilde{F}]\!=\!\mathrm{vec}^{-1}(\mu_{f}),~\mathbb{E}\left[L_{P^{-1}}\right]\!=\!L_{\Sigma_{p}}\mathsf{D}$ (using (8)) with $\Sigma_{p}\!=\!L_{\Sigma_{p}}L_{\Sigma_{p}}^{\top}$ and $\mathsf{D}$ a diagonal matrix with entries
$\mathsf{D}_{ii}={\mathbb{E}[c_{i}]}=\sqrt{2}\frac{\Gamma\left(\frac{{n}-i+2}{2}\right)}{\Gamma\left(\frac{{n}-i+1}{2}\right)}.$
@@
$\mathbb{E}[Q]=\begin{cases}Q,&\text{(WSN)},\\ k_{q}\Sigma_{q},&\text{(}Q\text{-WSN)},\\ \frac{k_{\alpha}\theta_{\alpha}}{k_{p}-n-1}\Sigma^{-1}_{P},&\text{(}\alpha P\text{-WSN},k_{p}>n+1\text{)}.\end{cases}$
The expectation formulas in (9) and the expression for the eigenvalues of $A$ reveal useful insights into how one may choose the parameters for the proposed distributions. Recall that the real part of any eigenvalue $\lambda$ of $A$ in (6) depends on $Q+\tilde{F}^{\top}\tilde{F}$, while the imaginary part is influenced by the skew-symmetric matrix $S$. For instance, when adopting the WNS distribution, one fixes $Q\in\mathbb{S}^{n}_{>0}$ and draws samples for $P^{-1}$, $\tilde{F}$, and $S$. This approach ties the mean-squared decay properties (through $Q$) to a known baseline while permitting flexibility in both the real part of the eigenvalues (through the distribution of $P$ and $\tilde{F}$) and their oscillatory behavior (through $S$). Suppose one wants to quantify the uncertainty in the mean-squared decay properties with the WNS. In that case, it is possible by either modifying the uncertainty in $P^{-1}$ through $\Sigma_{p}$ and $k_{p}$ or modifying the uncertainty in $\tilde{F}$ through $\Sigma_{f}$. However, this will also alter the uncertainty in the drift part’s oscillatory behavior or the diffusion part’s uncertainty. In contrast, the $Q$-WNS distribution treats $Q$ itself as random, thereby enabling us to quantify the uncertainty in mean-square decay properties without affecting the uncertainty in the oscillatory properties or the diffusion term.